1. How much heat is needed to raise the temperature of 3 kg of water from 20oC to 40oC? Give answer in kJ (1kJ = 1000 J) (round answer to 3 sig fig)


2. How much heat is needed to raise the temperature of 3 kg of ice from -40oC to -20oC? (round answer to 3 sig fig)


Specific heat of water: C = 4180 J/kgK
Specific heat of ice: C = 2090 J/kgK
Specific heat of ethanol: C= 2440 J/kgK
Answer #1

Heat required to raise the temperarure of the body is given as:

  Q=ms(T_{2}-T_{1})

here,

m = mass of the substance

s = specific heat of the substance

T_{1}= initial~temperature~of ~the~body~(before~heat~is ~given)T_{2}= final~temperature~of ~the~body(after~heat~is ~given)

1) Given:

  m = 3 kg

s = 4180 J /kg K

  T_{1}= 20^{\circ }C

T_{2}= 40^{\circ }C

So,

Heat required = 3* 4180 (40-20)

= 250,800 J

=250.800 kJ

heat required to raise the temperature of water is =250.800 kJ

......

2).  Given: m = 3kg

s = 2090 J/kg K

  T_{1}=- 40^{\circ }C

  T_{2}=- 20^{\circ }C

So, heat required = 3* 2090 (-20-(-40))

= 125,400 J

= 125.400 kJ

heat required to raise the temperature of ice is

125.400 kJ